4. Without doing any calculations, predict whether the observed ∆T would increase, decrease or remain the same if the following procedural changes were made. Justify your predictions:

a. 50.0 mL of 2.0 mol L NaOH mixed with 50.0 mL of 2.0 mol L HCl.

b. 25.0 mL of 2.0 mol L NaOH + 50.0 mL of water mixed with 25.0 mL of 2.0 mol L -1 -1 HCl.

5. Use tabulated rxn standard molar enthalpies of formation (from the textbook) to calculate ∆H 0 for the following (unbalanced) reactions:

a. N2O4(g) + H2 (g) ˇ N2 (g) + H2O(g)

b. H2S(g ) + O2 (g) ˇ SO2 (g) + H2O(g)

c. Fe2O3(s) + HCl(g) ˇ FeCl3 (s) + H2O(g)

6. One of the cleanest burning components of gasoline is 2,3,4-trimethylpentane which has the chemical formula C8H8.

a. Write a balanced chemical reaction for the complete combustion of 2,3,4- 2 2 trimethylpentane to CO (g) and H O(g) such that the stoichiometric coefficient for C8 1 H 8 is 1.

b. When 50.0 g of 2,3,4-trimethylpentane are completely burned at 298.15 K and 1 bar of pressure, 2.218 x 10 kJ are released. Calculate the ∆H of 2,3,4- 3 0 trimethylpentane.
Answer #1

Ans 4

Part a

50.0 mL of 2.0 mol/L NaOH mixed with 50.0 mL of 2.0 mol/L HCl

The balanced reaction

NaOH + HCl = NaCl + H2O

NaOH is strong base

HCl is strong acid

When strong acid mixes with strong base, heat is released in the reaction i.e. Exothermic reaction

Temperature before reaction (T1) < Temperature after reaction (T2)

(T2 - T1) > 0

∆T > 0

∆T would increase

Part b

b. 25.0 mL of 2.0 mol/L NaOH + 50.0 mL of water mixed with 25.0 mL of 2.0 mol/L HCl.

The balanced reaction

NaOH + HCl = NaCl + H2O

NaOH is strong base

HCl is strong acid

Volume of acid and base = 25 + 25 = 50 mL

Volume of water = 50 mL

When strong acid mixes with strong base, heat is released in the reaction i.e. Exothermic reaction but 50 mL water is also added to neutralize the heat.

Temperature before reaction (T1) = Temperature after reaction (T2)

(T2 - T1) = 0

∆T = 0

∆T would remain same